# Two-level Nonlinear Preconditioning Methods for Flood Models Posed on Perforated Domains

### v1

@@ 3 Multiscale Coarse space
$V_{H}=\{\,v\in H_{\Delta}^{1}(\Omega)\,\,|\,\,v|_{\Gamma}\in V_{H}^{\Gamma}\},$
where
@@ 4 Linear Multi-domain Solution Methods
Generally, the number of iterations in (10) and (11) grows with the number of subdomains in one direction. For example, this increase is linear for the stationary iteration (10). To combat this, a coarse level is needed to aid in global communication between all subdomains. Consider the coarse matrix $R_{H}$ defined in Section 3; with this, the two-level preconditioned Krlov method method is given by
@@ 5 Nonlinear Multi-domain Solution Methods
To solve nonlinear equations, Newton’s method is commonly employed. For exact Newton’s method, one needs the the residual $F(\mathbf{u})$ as well as the Jacobian $\nabla F(\mathbf{u})$. The exact Newton’s method is described in Algorithm 1.
@@
for $k=0,\ldots,$ to convergence do
@@
Often when implementing Newton’s method, a backtracking line search is implemented to control the step size of Newton’s method. This often aids in convergence for certain problems as the line search prevents “overshooting” in the Newton step direction. Newton’s method with backtracking line search is given in Procedure 1.
@@
for $k=0,\ldots,$ to convergence do
@@
When a linear system is very large, a direct solve is often too expensive and not recommended. Therefore, it is common to solve the linear system in Algorithm 1 via a preconditioned Krylov solver. We proposed in [4] and recalled in Section 3 a coarse space for the two-level RAS preconditioner applied to the Poisson equation on perforated domains. It turns out that the same coarse space can be successfully employed within the two-level RAS preconditoner (13) for the linear system in Algorithm 1. The numerical evidences regarding the efficiency of such a two-level solver are presented in Section 6.
@@
An NRAS iteration is computed by first solving a nonlinear subsystem corresponding to $\Omega_{j}^{\prime}$ then gluing these local subdomain updates together. Let $R_{j}$ be the boolean restriction matrices corresponding to the degrees of freedom of $\Omega_{j}^{\prime}$; we introduce the “subdomain update” function $G_{j}$, such that, for all $\mathbf{u}$, $G_{j}(\mathbf{u})$ is the solution of
@@
The subdomain updates $G_{j}(\mathbf{u})$ correspond to solving the problem on $\Omega_{j}^{\prime}$ with Dirichlet boundary conditions. The local problems (15) are solved via Newton’s method; as the subproblems are small, each Newton linear system is solved using a direct solver. Then, the local solutions $G_{j}(\mathbf{u})$ are glued together, giving the global update function
@@
We remark that, in principle, as in the linear case, we can use this nonlinear RAS method as a fixed point iteration, giving
@@
As a quick remark, we comment on the connection between the nonlinear RAS fixed point iteration (17) and the linear RAS fixed point iteration (10). Consider the local solutions computed from (15). For linear problems such that $F(\mathbf{u})=A\mathbf{u}-\mathbf{f}$, we have
@@
resulting in one-level RASPEN method [13] detailed in Algorithm 2. It was shown in [13] that the Jacobian of $\mathcal{F}(\mathbf{u})$ can be computed directly.
@@
for $k=0,\ldots,$ to convergence do
@@
We now describe the computation of the Jacobian $\nabla\mathcal{F}(\mathbf{u}^{k})$ for Algorithm 2. In view of (16), we have
@@
$\nabla\mathcal{F}(\mathbf{u}^{k})$ $=\nabla\left(\mathbf{u}^{k}-\sum_{j}R_{j}^{T}D_{j}G_{j}(\mathbf{u}^{k})\right)$ (20)
$=I-\sum_{j}R_{j}^{T}D_{j}\nabla G_{j}(\mathbf{u}^{k})$
Therefore, to compute the Jacobian $\nabla[\mathcal{F}(\mathbf{u}^{k})]$, we must compute $\nabla G_{j}(\mathbf{u}^{k})$. To compute $\nabla G_{j}(\mathbf{u}^{k})$, recall (15). Taking the derivative of this equation and solving for $\nabla G_{j}(\mathbf{u}^{k})$ gives
@@
where we have used the partition of unity property $I=\sum_{j}R_{j}^{T}D_{j}R_{j}$. We note that the inverse $[R_{j}\nabla F(\mathbf{v}_{j}^{k})R_{j}^{T}]^{-1}$ can be dense matrix. Therefore, we set up $\nabla\mathcal{F}(\mathbf{u}^{k})$ as a linear operator, allowing us to compute the matrix-vector product $\nabla\mathcal{F}(\mathbf{u}^{k})x^{k}$ for any vector $x^{k}$ without explicitly computing the matrix $\nabla\mathcal{F}(\mathbf{u}^{k})$.
@@
We choose a multiplicative two-level RASPEN correction provided in [19], which allows the coarse correction to be done algebraically. The authors of [15] provide numerical experiments on multiple different choices of two-level methods, including adding the coarse correction before, after, and before and after the nonlinear RAS iteration. The multiplicative two-level RASPEN method we have chosen is given by Algorithm 3. This method solves the nonlinearly preconditoned system
@@
for $k=0,\ldots,$ to convergence do
@@
Next, we consider a Two-step method proposed in [9] under the name NKS-RAS; see also [5]. This method alternates between an NRAS iteration and a global Newton step at each outer iteration. This method leads to fast convergence (typically quadratic convergence is expected), but unlike Algorithms 2 and 3, can not be framed as a nonlinear preconditioning method. The Two-step method is given in Algorithm 4. The main feature of the Two-step method is the ease of its implementation. Compared to the standard Newton’s method, it simply requires an additional NRAS step; moreover, compared to one and two-level RASPEN, the two-step method does not involve any dense linear systems, and therefore, in principle, may be implemented based on a direct linear solver. Here, the linearized system of Algorithm 4 is solved using the same preconditioned GMRES method as in Algorithm 1.
@@
for $k=0,\ldots,$ to convergence do
@@
for $k=1,2,\ldots,$ to convergence do
@@
$\min_{\alpha^{k}=(\alpha_{0}^{k},\ldots,\alpha_{m_{k}}^{k})^{T}}||f_{k}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha^{k}}||_{2}\quad\text{such that}\quad\sum_{i=0}^{m_{k}}\alpha_{i}^{k}=1;$
@@ 5.1 Complexities/Costs of Fine-scale Methods
We now provide an overview of the complexity of each method. To quantify the cost of each algorithm, we recall that at each time step, each outer iteration consists of the following for each method:
@@
    Newton: 1 global Newton update (1 global two-level linearly preconditioned Krylov solve);
@@
    Two-Step: 1 NRAS Update + 1 global Newton update (1 global two-level linearly preconditioned Krylov solve);
@@
    One-level RASPEN: 1 NRAS update + 1 global Newton update (1 one-level nonlinearly preconditoned Krylov solve);
@@
    Two-level RASPEN: 1 NRAS update + 1 coarse nonlinear solve + 1 global Newton update (1 two-level nonlinearly preconditioned Krylov solve);
@@
    Anderson: 1 NRAS update + 1 coarse linear solve + 1 small least-squares problem.
Computing the NRAS update (16) requires solving a family of local nonlinear systems associated to the subdomains. The local nonlinear systems are computed by Newton’s method, with a linear system solved at each iteration. We denote by $N_{it,loc}$ the required number of local Newton iterations averaged over all subdomains. Due to the small size of the subdomains, the linearized system at each Newton iteration is computed via a direct sparse linear solve. We denote by $C(LU_{loc})$ a typical cost of the local LU decompositions on each subdomain, assumed to be approximately the same for every subdomain. The complexity of the NRAS iteration can be expressed by
$C(\text{NRAS})$ $\approx N\cdot C(LU_{loc})N_{it,loc},$ (27)
Here, (27) assumes that the local computations are not done in parallel.
With the exception of Anderson’s method, all methods involve a global linear system which is solved via GMRES. The complexity of the GMRES method is given by
$N\cdot C(LU_{loc})+N_{\text{GM}}C(\text{GMRES}),$
where $N\cdot C(LU_{loc})$ denotes the one-time cost of assembling the preconditioner and $C(\text{GMRES})$ denotes the cost of each GMRES iteration, which is generally involves a sparse matrix-vector multiplication. We note that in the RASPEN variants, the LU decompositions can be reused from the earlier NRAS iteration; this would slightly reduce the cost further. We denote by $C(\text{coarse})$ the cost of the sparse linear solve (25). Therefore, $N_{\text{c}}C(\text{coarse})$ denotes the total cost of the coarse nonlinear problems in the two-level RASPEN algorithm (Algorithm 3), where $N_{\text{c}}$ denotes the number of coarse linear solves.
Furthermore, we quantify the following costs:
@@
where the LU decomposition of some sparse $m\times m$ matrix is supposed to be $O(m^{3/2})$ [16]. Finally, for each outer iteration, each method has the following cost:
$C$(Newton) $\approx$            $N\cdot O(N_{loc}^{3/2})+N_{\text{GM}}O(N_{\Omega})$;
$C$(Two-Step) $\approx$ $N\cdot O(N_{loc}^{3/2})N_{it,loc}+N\cdot O(N_{loc}^{3/2})+N_{\text{GM}}O(N_{\Omega})$;
$C$(One-RASPEN) $\approx$ $N\cdot O(N_{loc}^{3/2})N_{it,loc}+N\cdot O(N_{loc}^{3/2})+N_{\text{GM}}O(N_{\Omega})$;
$C$(Two-RASPEN) $\approx$ $N\cdot O(N_{loc}^{3/2})N_{it,loc}+N\cdot O(N_{loc}^{3/2})+N_{\text{GM}}O(N_{\Omega})+N_{\text{c}}O(N_{\cal V}^{3/2})$;
$C$(Anderson) $\approx$ $N\cdot O(N_{loc}^{3/2})N_{it,loc}+\qquad\,\,\,\qquad\qquad\qquad O(N_{\Omega})+\,\quad O(N_{\cal V}^{3/2})$;
where the $O(N_{\Omega})$ term arises as the dominant term in the least-squares problem of Anderson’s method. Clearly, the values $N_{it,loc},N_{\text{GM}},N_{\text{c}}$ and the total number of outer iterations can not be known a priori. In Section 6, we record the number of GMRES iterations $N_{\text{GM}}$, the number of local linear NRAS iterations averaged over all subdomains $N_{it,loc}$, and the total number of coarse linear iterations $N_{\text{c}}$, as well as the total number of outer iterations.

### v2

@@ 3 Multiscale Coarse space
$V_{H}=\{\,v\in H_{\Delta}^{1}(\Omega)\,\,|\,\,v|_{\Gamma}\in V_{H}^{\Gamma}\}.$
@@ 4 Linear Multi-domain Solution Methods
Generally, the number of iterations in (10) and (11) grows with the number of subdomains in one direction. For example, this increase is linear for the stationary iteration (10). To combat this, a coarse level is needed to aid in global communication between all subdomains. Consider the coarse matrix $R_{H}$ defined in Section 3; with this, the two-level preconditioned Krylov method is given by
@@ 5 Nonlinear Multi-domain Solution Methods
To solve nonlinear equations, Newton’s method is commonly employed. For exact Newton’s method, one needs the residual $F(\mathbf{u})$ as well as the Jacobian $\nabla F(\mathbf{u})$. The exact Newton’s method is described in Algorithm 1.
@@
for outer iteration $k=0,\ldots,$ to convergence do
@@
Often when implementing Newton’s method, a backtracking line search is implemented to control the step size of Newton’s method. This often aids in convergence for certain problems, as the line search prevents “overshooting” in the Newton step direction. Newton’s method with backtracking line search is given in Procedure 1.
@@
for outer iteration $k=0,\ldots,$ to convergence do
@@
When a linear system is very large, a direct solve is often too expensive and not recommended. Therefore, it is common to solve the linear system in Algorithm 1 via a preconditioned Krylov solver. We proposed in [4] and recalled in Section 3 a coarse space for the two-level RAS preconditioner applied to the Poisson equation on perforated domains. It turns out that the same coarse space can be successfully employed within the two-level RAS preconditioner (13) for the linear system in Algorithm 1. The numerical evidences regarding the efficiency of such a two-level solver are presented in Section 6.
@@
Definition 5.1 (An NRAS iteration).
Let $R_{j}$ be the boolean restriction matrices corresponding to the degrees of freedom of $\Omega_{j}^{\prime}$. Let the “subdomain update” function $G_{j}$, such that, for all $\mathbf{u}$, $G_{j}(\mathbf{u})$ is the solution of the local nonlinear system
@@
An NRAS iteration is computed by first solving each (15) then glue local contributions:
@@
We can use this method as a fixed point iteration, computing iteratively the solution $\bf{u}^{k}$ starting from an initial guess $\bf{u}^{0}$:
@@
Note that the subdomain updates $G_{j}(\mathbf{u})$ correspond to solving the local problem on $\Omega_{j}^{\prime}$ with Dirichlet boundary conditions. These problems (15) are solved via Newton’s method; as the subproblems are small, each Newton linear system is solved using a direct solver.
Remark 5.1 (NRAS is RAS when applied to a linear problem).
We comment here on the connection between the nonlinear RAS fixed point iteration (17) and the linear RAS fixed point iteration (10). Consider the local solutions computed from (15). For linear problems such that $F(\mathbf{u})=A\mathbf{u}-\mathbf{f}$, we have
@@
resulting in the One-level RASPEN method [15] detailed in Algorithm 2. It was shown in [15] that the Jacobian of $\mathcal{F}(\mathbf{u})$ can be computed directly.
@@
for outer iteration $k=0,\ldots,$ to convergence do
@@
Remark 5.2 (Exact computation of the Jacobian).
Unlike in the case of ASPIN, for RASPEN the Jacobian is computed exactly. We describe here the computation of the Jacobian $\nabla\mathcal{F}(\mathbf{u}^{k})$ for Algorithm 2. From (16), we have
@@
$\nabla\mathcal{F}(\mathbf{u}^{k})$ $=\nabla\left(\mathbf{u}^{k}-\sum_{j}R_{j}^{T}D_{j}G_{j}(\mathbf{u}^{k})\right)=I-\sum_{j}R_{j}^{T}D_{j}\nabla G_{j}(\mathbf{u}^{k}).$ (20)
To evaluate this, we must compute $\nabla G_{j}(\mathbf{u}^{k})$ by the gradient of (15):
@@
where we have used the partition of unity property $I=\sum_{j}R_{j}^{T}D_{j}R_{j}$. We note that the matrix $[R_{j}\nabla F(\mathbf{v}_{j}^{k})R_{j}^{T}]^{-1}$ can be dense and does not need to be directly assembled. Therefore, we set up $\nabla\mathcal{F}(\mathbf{u}^{k})$ as a linear operator, allowing us to compute the matrix-vector product $\nabla\mathcal{F}(\mathbf{u}^{k})x^{k}$ for any vector $x^{k}$ without explicitly computing the matrix $\nabla\mathcal{F}(\mathbf{u}^{k})$.
@@
We choose a Two-level RASPEN correction provided in Algorithm 11 of [21], which allows the coarse correction to be done algebraically. The coarse correction is chosen to be computed after the local nonlinear solves. The authors of [17] provide numerical experiments on multiple different choices of two-level methods, including adding the coarse correction before, after, and before and after the nonlinear RAS iteration. The multiplicative Two-level RASPEN method we have chosen is given by Algorithm 3. This method solves the nonlinearly preconditioned system
@@
for outer iteration $k=0,\ldots,$ to convergence do
@@
Next, we consider a Two-step method proposed in [9] under the name NKS-RAS; see also [5]. This method alternates between an NRAS iteration and a global Newton step at each outer iteration. This method leads to fast convergence (typically quadratic convergence is expected), but unlike Algorithms 2 and 3, can not be framed as a nonlinear preconditioning method. The Two-step method is given in Algorithm 4. The main feature of the Two-step method is the ease of its implementation. Compared to the standard Newton’s method, it simply requires an additional NRAS step; moreover, compared to the One and Two-level RASPEN methods, the Two-step method does not involve any dense linear systems, and therefore, in principle, may be implemented based on a direct linear solver. Here, the linearized system of Algorithm 4 is solved using the same preconditioned GMRES method as in Algorithm 1.
@@
for outer iteration $k=0,\ldots,$ to convergence do
@@
for outer iteration $k=1,2,\ldots,$ to convergence do
@@
$\min_{\alpha^{k}=(\alpha_{0}^{k},\ldots,\alpha_{m_{k}}^{k})^{T}}||f_{k}\alpha^{k}||_{2}\quad\text{such that}\quad\sum_{i=0}^{m_{k}}\alpha_{i}^{k}=1;$
@@ 5.1 Complexities/Costs of Fine-scale Methods
We now provide an overview of the complexity of each method. Recall that all proposed methods besides Newton’s method involve an NRAS update. Computing the NRAS update (16) requires solving a family of local nonlinear systems associated with the subdomains. The local nonlinear systems are computed by Newton’s method, with a linear system solved at each iteration. We denote by $N_{it,loc}$ the required number of local Newton iterations averaged over all subdomains. Due to the small size of the subdomains, the linearized system at each Newton iteration is computed via a direct sparse linear solve. This number might vary according to the method used, and we use an upper index to distinguish different case scenarios. We denote by $C(LU_{loc})$ a typical cost of the local LU decompositions on each subdomain, assumed to be approximately the same for every subdomain. Additionally, we denote by $C(\text{local assembly})$ the cost of assembling the local residual and Jacobian vectors, which will be done for all subdomains at each local iteration. The complexity of the NRAS iteration at each outer iteration can be expressed by:
$C(\text{NRAS})$ $=N\cdot N_{it,loc}\left(C(LU_{loc})+C(\text{local assembly})\right),$ (27)
where $N$ is the number of subdomains. Here, (27) assumes that the local computations are not done in parallel. With a parallel implementation, instead of the NRAS cost being summed over all subdomains, the cost would be dominated by the subdomain associated to the maximum number of local linear solves.
With the exception of Anderson’s method, all methods involve a global linear system which is solved via GMRES at each iteration. The complexity of the GMRES method for the Two-step and Newton methods at each outer iteration is given by the following:
$N\cdot C(LU_{loc})+C(LU_{coarse})+C(\text{assembly})+N_{\text{GM}}\cdot C(\text{GMRES}),$ (28)
where $N\cdot C(LU_{loc})+C(LU_{coarse})$ denotes the cost of assembling the preconditioner at each outer iteration, $C(\text{assembly})$ denotes the cost of assembling the global residual and Jacobian vectors/matrices $F({\bf u}),\nabla F({\bf u})$, and $C(\text{GMRES})$ denotes the cost of each GMRES iteration, which generally involves a sparse matrix-vector multiplication. We denote by $N_{\text{GM}}$ the average number of GMRES iterations per outer iteration. That is, $N_{\text{GM}}$ denotes the average number of GMRES iterations until convergence of the preconditioned linear systems occurring at each outer iteration, which will depend on the chosen solution strategy. We use the upper index to $N_{\text{GM}}^{i}$ to illustrate that this number can be different and decrease with the addition of an accelerator like a coarse space.
We note that in the RASPEN variants, the LU decompositions and assemblies for the nonlinearly preconditioned system can be reused from the earlier NRAS iteration, reducing the cost of the GMRES method compared to (28). That is, the complexity of the GMRES method for the One and Two-level RASPEN methods at each outer iteration is given by the following:
$N_{\text{GM}}\cdot C(\text{GMRES}),$
where this cost is equal to (28) for the Newton and Two-step methods. We denote by $C(\text{coarse})$ the cost of the sparse linear solve used for the coarse nonlinear problem in the Two-level RASPEN algorithm. Therefore, $N_{\text{c}}\cdot C(\text{coarse})$ denotes the total cost of the coarse nonlinear problems per outer iteration, where $N_{\text{c}}$ denotes the average number of coarse linear solves per outer iteration.
With this, the costs for each method are summarized to be
@@
    Newton: $N\cdot C(LU_{loc})+C(LU_{coarse})+C(\text{assembly})+N_{\text{GM}}\cdot C(\text{GMRES})$;
@@
    Two-Step: $C(\text{NRAS})+N\cdot C(LU_{loc})+C(LU_{coarse})+C(\text{assembly})+N_{\text{GM}}\cdot C(\text{GMRES})$;
@@
    One-level RASPEN: $C(\text{NRAS})+N_{\text{GM}}\cdot C(\text{GMRES})$;
@@
    Two-level RASPEN: $C(\text{NRAS})+N_{\text{GM}}\cdot C(\text{GMRES})+N_{c}\cdot C(\text{coarse})$;
@@
    Anderson: $C(\text{NRAS})+C(\text{assembly})+C(\text{coarse})$.
Furthermore, we quantify the following costs as follows:
@@
    $C(LU_{coarse})=O(N_{\cal V}^{3/2})$;
  • 3.
    $C(\text{local assembly})=O(N_{loc})$;
  • 4.
    $C(\text{assembly})=O(N_{\Omega})$;
  • 5.
@@
where the LU decomposition of some sparse $m\times m$ matrix is supposed to be $O(m^{3/2})$ [18].
We denote the total number of outer iterations by $N_{k}$; this number may vary according to the method used, and we can use an upper index to distinguish between methods. Finally, the total cost over all outer iterations is given by:
$C(\text{Newton})$ $=N_{k}^{0}\Bigl{(}$ $N$$\times$ $O(N_{loc}^{3/2})$ $+$ $(N_{\text{GM}}^{0}+1)$$\times$ $O(N_{\Omega})$ $+$ $O(N_{\cal V}^{3/2})\Bigr{)}$;
[…]
$C(\text{Anderson})$ $=N_{k}^{4}\Bigl{(}$ $N\times N_{it,loc}^{4}$$\times$ $O(N_{loc}^{3/2})$ $+$ $O(N_{\Omega})$ $+$ $O(N_{\cal V}^{3/2})\Bigr{)}$;
where the $O(N_{\Omega})$ term arises as the dominant term in the least-squares problem of Anderson’s method and $C(\text{assembly})$ is included in Anderson’s method as well. Additionally, we have simplified the cost of the local assembly as $N\cdot N_{it,loc}\cdot O(N_{loc})\approx N_{it,loc}\,O(N_{\Omega})$ from the definition of $N_{loc}$.
We remark that for a time-dependent problem such as (1), there will be an additional $N_{T}$ term multiplied into the cost of each method, where $N_{T}$ denotes the number of time steps. In this case, $N_{k}^{i}$ would denote the average number of outer iterations per time step. As shown, we can estimate the complexity of individual components of the methods such as the cost of assembly, the cost of a sparse linear solve, and the cost of a matrix-vector product. Additionally, we report the iteration counts such as $N_{k}^{i}$, $N_{\text{GM}}^{i},$ $N_{it,loc}^{i}$, and $N_{c}$ in Section 6.
For the Two-step, Newton, and Two-level RASPEN methods, the GMRES method occurring at each outer iteration will be preconditioned with some two-level DD preconditioner. Therefore, we anticipate that $N_{GM}^{0},N_{GM}^{1}$, and $N_{GM}^{3}$ will be approximately equal, with $N_{GM}^{2}$ for the One-level RASPEN method growing as we increase $N$. We also anticipate that $N_{it,loc}^{i}$ will be consistent between all methods which involve an NRAS solve (each method besides Newton). Therefore, the cost of the Two-step and Two-level RASPEN methods can be directly compared by comparing the the outer iterations $N_{k}$, where $N_{\cal V}<<N_{\Omega}$ such that the cost of a coarse linear solve will be significantly lower than a fine linear solve. For reference, we will still record the number of coarse linear iterations $N_{c}$ for the Two-level RASPEN to ensure it is not dominating the cost.
For Newton’s method, it is clear that the cost of each iteration is lower than the Two-step and RASPEN methods, as there is no NRAS solve involved. However, the main goal of the proposed methods is to substantially reduce the number of outer iterations to make up for the increased complexity within each outer iteration. That is, we expect $N_{k}$ to be reduced for the other methods, allowing for a potential gain in performance. With regard to Anderson’s method, it appears to have quite a low cost at each iteration, with one NRAS solve at each iteration and the cost of the least-squares problem dominated by one $O(N_{\Omega})$ term. However, we do expect a higher value of outer iterations for this method $N_{k}^{4}$ when compared to the Two-step and RASPEN variants. We note that while it is difficult to directly compare an outer iteration of Anderson to the other methods, it is still an interesting alternative that is worth exploring in this article.
With this, our main concern in terms of complexity lies in the term $N_{k}^{i}N_{\text{GM}}^{i}O(N_{\Omega})$, representing the total number of “inner” GMRES iterations. However, when applicable, we also wish to record the term $N_{k}^{i}N_{it,loc}^{i}$, representing the total number of sparse local linear solves on a given subdomain. In Section 6, we record the number of outer iterations $N_{k}^{i}$, the average number of GMRES iterations per outer iteration $N_{\text{GM}}^{i},$ the cumulative number of GMRES iterations $N_{k}^{i}N_{\text{GM}}^{i}$, as well as the average and cumulative (over all outer iterations) local linear solves per subdomain, namely $N_{it,loc}^{i}$ and $N_{k}^{i}N_{it,loc}^{i}$.